OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2)

Parth (parthkohli):

Distribute first

OpenStudy (anonymous):

5x +40 =7x-14 => 5x-7x = -14-40 => -2x =-54 => x = 54/2 = 27

mathslover (mathslover):

\[a(b+c) =a(b)+a(c) = ab+ac\] \[5(x+8)=7(x-2)\] \[5x+5(8) = 7(x)-7(2)\] \[5x+40 = 7x-14\] \[5x-7x = -14-40\] \[-2x = -54\] \[x=54/2 = 27\]